# Federated Learning Clients Clustering with Adaptation to Data Drifts

### v1

@@ 4. Fielding Overview
Figure 3. Clustering overview of Fielding. Clients send distribution vectors to the coordinator; the coordinator moves drifted clients to the closest cluster; global re-clustering is triggered if any cluster center shifts by a distance larger than $\tau$.
To improve clustered FL performance, we develop Fielding, a clustered FL framework that efficiently reassigns clients after data drifts. As explained in Section 3, Fielding has three key design decisions: (1) reassign all drifted clients instead of only selected clients, (2) combine per-client adjustment and global re-clustering, and (3) cluster clients based on label distribution vectors.
4.1. System Overview
As discussed in Section 3.1, neither per-client adjustment nor global re-clustering works well across all drift scenarios. Therefore, we propose a combined approach that leverages both methods. Fielding employs a threshold $\tau$ to determine whether global re-clustering is necessary. When clients report data drifts, we first try re-clustering drifted clients individually by moving them to the cluster whose center is the closest to their new distribution vector. To ensure determinism across different orders of client movements, we do not update the cluster center throughout this per-client adjustment process. After individual movements, we recalculate the cluster centers and trigger global re-clustering over all clients only when any center moves by a distance larger than $\tau$. In our implementation, we set $\tau$ as one-third of the average distance between a pair of cluster centers, which is a value that works empirically well across evaluation settings.
Figure 3 illustrates the clustering mechanism of Fielding. When clients register with the centralized coordinator, they send their local data distribution vector, either during initial registration or following a recent data drift. Upon receiving this vector, the coordinator updates the client’s metadata and assigns the client to the nearest cluster. Before selecting participants for the upcoming training round, the coordinator finds the distance each cluster center has shifted. It then initiates a global clustering process if any cluster exceeds the threshold $\tau$. Otherwise, it finalizes the updated cluster membership.
Our clustering approach incurs additional communication costs, as clients are expected to upload their data distribution vector. However, this vector is small (4 bytes×$\times N$ where $N$ is the number of labels), and clients only need to send it when they first join or experience data drift. In Section 4.3, we prove the model convergence guarantees with label-based clustering.
4.2. Algorithmic Framework
@@
1 Partition clients into clusters using Algorithm 3
@@ 2 Initialize cluster models $\mathbf{c}^{(1)},\ldots,\mathbf{c}^{(K)}$
4       Handle data drifts using Algorithm 2
@@ 15                  $\mathbf{c}^{(k)}\leftarrow\operatorname*{avg}_{i\in S}\mathbf{x}_{i}$
Algorithm 1 Clustered Federated Learning with data drift
@@ 7             Let $\mathbf{x}_{i}$ be the model corresponding to the $i$’th client’s cluster
8      Let $C_{1},\ldots,C_{K}$ be clusters returned by Algorithm 3
@@ 10             $\mathbf{c}^{{(k)}}\leftarrow\operatorname*{avg}_{i\in C_{k}}\mathbf{x}_{i}$
Algorithm 2 Handling data drifts
1 Define the distance between clients as the $\ell_{1}$-distance between their label histograms
@@ 2 Choose $K$ with the largest silhouette score
In this section, we describe our algorithm – presented in Algorithm 1 – in detail. Initially, Fielding clusters the clients using the $k$-means algorithm (Algorithm 3). Importantly, we measure distance between data distributions, which is computed as the $\ell_{1}$-distance between their histograms: that is, assuming we have $l$ labels, the distance between clients with label distribution $p_{1},\ldots,p_{l}$ and $q_{1},\ldots,q_{l}$ is defined as $\sum_{i=1}^{l}|p_{i}-q_{i}|$.
We assume that the algorithm is executed for $T$ global iterations, and during each iteration, the label distribution of each client might change, for instance by adding or removing data points. The first step of Fielding is to handle the data drifts (Algorithm 2). Our goal is, by the end of each iteration, to maintain a good model for each cluster, with respect to its clients’ data. In this step, each drifted client is assigned to the closest cluster. If this causes a significant change in cluster centers – namely, if there exists a cluster moving by more than $\tau/3$, where $\tau$ is the average distance between cluster centers – we use Algorithm 3 to recluster clients from scratch. This condition avoids frequent reclustering, which might require excessive resources and can adversely change the clients’ losses. Importantly, we recluster all the clients, not just available clients, which, as we show in Section 2, improves the accuracy.
For the global reclustering, we choose the number of clusters that gives the highest silhouette score. After reclustering, we compute new cluster models: for each client $i$, let $\mathbf{x}_{i}$ be its old cluster model, and then for each new cluster $C_{k}$, we define its model as an average of $\mathbf{x}_{i}$ for $i\in C_{k}$. That is, we average the old client models.
After the clusters are computed, we train the cluster models for $R$ rounds. During each round, we sample the clients from the cluster and run $L$ local iterations of gradient descent. After the local iterations, we set the cluster model as the average of the models of all participating clients.
4.3. Convergence
In Appendix A, we analyze the performance of our framework by bounding the average of clients’ loss functions at every iteration. Formally, for every cluster $C_{k}$, let $\mathbf{c}^{(k)}$ be the model corresponding to the cluster by the end of the iteration, and let $\mathbf{c}^{(k,*)}$ be the optimal model for this cluster, that is, the minimizer of $\sum_{i\in C_{k}}f^{(i)}_{t}$. Then, we bound the average loss of the clients with respect to the best possible cluster models:
(1) $\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in C_{k}}(f^{(i)}_{t}(\mathbf{c}^{(k)})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}))$
Our analysis has two ingredients, corresponding to two main stages in our algorithm: gradient descent steps and reclustering. For the gradient descent part, we use the standard convergence analysis of SGD, which for strongly convex function $f$ shows that for stochastic gradient descent update the following holds:
$\mathbb{E}\left[f(\mathbf{x}_{T})-f(\mathbf{x}^{*})\right]$ $\leq\left(1-\eta\mu\right)^{T}(f(\mathbf{x}_{0})-f(\mathbf{x}^{*}))+\frac{L\eta}{2\mu}\sigma^{2}$
The right-hand side has two terms, where the first term decays exponentially, and the second term corresponds to the total contribution of the stochastic noise. In our analysis, we similarly have the $(f(\mathbf{x}_{0})-f(\mathbf{x}^{*}))\left(1-\eta\mu\right)^{T}$ term which decays exponentially, and a non-vanishing term corresponding to the noise coming from the stochastic noise and the client sampling.
The major challenge in the analysis is bounding the effect of data drifts and reclustering. Data drifts change client datasets, hence changing their local objective functions. On the other hand, while reclustering is useful in the long run, its immediate effect on the objective (1) can potentially be negative (see Figure 2(b)) due to the mixing of models from different clusters.
To address these issues, we introduce natural assumptions. First, we assume that the distance between client representations – which in our algorithm correspond to label distributions – translates into the difference between their objective functions. Second, we assume that the effect of data drift is bounded: that is, the representation $r_{i}$ of client $i$ changes by at most $\delta$ for each data drift. Finally, we assume that the clients are clusterable; that is, $K$ clusters exist, so representations of clients within each cluster are similar.
@@ 4.3. Convergence
$f^{(i)}_{t}$ – local objective for client $i$ at iteration $t$ – is $\mu$-strongly convex and $L$-smooth, and we have access to a stochastic oracle with variance $\sigma^{2}$. There exists clustering such that representations inside each cluster are $\Delta$-close, the data drift changes representations by at most $\delta$, and the difference between the objective functions is at most the distance between representations multiplied by $\theta$.
Under these assumptions, the data drift can affect each client representation – and hence objective 1 – by at most a fixed value. Moreover, by the clustering assumption, the clients within each cluster have similar objective function; moreover, this holds both before and after the reclustering. This allows us to bound the increase in the loss function due to reclustering. In Appendix A, we prove the following theorem.
Theorem 4.1.
@@
$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$
$\leq(1-\eta\mu)^{TR}\left(f_{0}(\mathbf{x}_{0})-f_{0}(\mathbf{x}^{*})\right)$
$+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
That is, objective (1) has two terms: an exponentially decaying term corresponding to the initial loss, and a non-vanishing term corresponding to stochastic noise, as well as loss due to clustering and data drift.
5. Implementation
We implemented Fielding in Python on top of FedScale (Lai et al., 2022), a state-of-the-art open-source FL engine. We follow FedScale’s design of having one centralized coordinator and client-specific executors. Figure 4 illustrates the end-to-end workflow of Fielding. At the beginning of each training round, 1 client executors register with the coordinator to participate and report their local data distribution optionally. The coordinator’s cluster manager maintains clients’ distribution records and 2 moves clients to the closest cluster when their reported distribution drifts. After moving drifted clients individually, the cluster manager measures center shift distances and either triggers global re-clustering when any cluster center shifts significantly or finalizes clusters membership. In the case of global clustering, we use $K$-means clustering and determine $K$ using the silhouette method (Rousseeuw, 1987). To stabilize new clusters and avoid creating poor clusters due to bad centroids initialization (Ahmed et al., 2020), we reuse previous cluster centers when the target $K$ value is no greater than the number of existing clusters. The initial model of each newly created cluster is set as the average of its clients’ previous cluster model (see Algorithm 2).
Figure 4. The system architecture and workflow of Fielding. Blue arrows represent communication between clients and the coordinator. Black arrows indicate the workflow among different components on the same machine.
6. Evaluation
We evaluate Fielding’s effectiveness by running four FL tasks. We also demonstrate Fielding’s compatibility with various client selection strategies and aggregation algorithms. Our end-to-end results highlight that
  • Fielding improves the final test accuracy by 1.9%-5.9% and is more stable than other clustered FL works when facing real-world data drifts.
  • Fielding works well with complementary FL optimizations and offers consistent accuracy improvements across different settings.
  • Fielding is robust to malicious clients who report corrupted distribution vectors and accelerates model convergence when clients are less heterogeneous.
Figure 5. Time to accuracy (TTA) comparison over four tasks.
@@ 6. Evaluation
Environment. We emulate large-scale FL training with four NVIDIA A100 GPUs. We use FedScale’s device datasets which provide realistic device computing and network capacity numbers while following the standardized training configurations in the original paper (Lai et al., 2022). Per-device computation time estimation is based on device computing speed, batch size, and number of local iterations. And we estimate the communication time as the volume of downloaded and uploaded data divided by the network bandwidth.
@@
Datasets and Models. We use a satellite image dataset (FMoW (Christie et al., 2018)) and two video datasets (Waymo Open (Sun et al., 2020) and Cityscapes (Cordts et al., 2016)) that include natural data drifts. For FMoW, we keep images taken after January 1, 2015, and create one client for each UTM zone. Two training rounds correspond to one day in this dataset. For Waymo Open and Cityscapes, we use video segments pre-processed by Ekya (Bhardwaj et al., 2022). We create one client per camera per Waymo Open segment (212 clients in total), and one client per 100 consecutive frames per Cityscapes segment (217 clients in total). As these two datasets have video frame IDs within segments but lack global timestamps, we sort the samples on each client by frame ID and then partition the frames into 10 intervals. We stream in one data interval every 20 training rounds. We train ResNet-18 (He et al., 2016) on these three real-world datasets.
@@
To demonstrate how Fielding behaves in highly dynamic settings, we construct an artificial trace using the FedScale Open Images (Kuznetsova et al., 2020) benchmark. We find the top 100 classes in terms of number of samples and keep clients whose local data labels include at least 10 of the top 100 classes (5078 clients in total). Each client randomly and independently partitions all local data labels into 10 buckets. Each bucket then contains all local samples of labels mapped to it. We stream in one bucket every 50 training rounds. We train ShuffleNet v2 (Ma et al., 2018) on Open Images. On all four datasets, clients retain samples they received over the last 100 rounds. The details of experiment settings and training parameters are included in Appendix B.
@@
Metrics. Our main evaluation metrics are Time-to-Accuracy (TTA) and final test accuracy. We define TTA as the training time required to achieve the target accuracy, which is the average test accuracy when training a single global model for all clients. We report the final average test accuracy across diverse settings to demonstrate the sensitivity and robustness of Fielding.
6.1. End-to-end Training Performance
Figure 5 shows that Fielding improves clients’ average test accuracy by 5.9%, 1.9%, and 2.1% on FMoW, Cityscapes, and Waymo Open respectively. When considering the global model’s final test accuracy as the target accuracy, Fielding offers a 1.27$\times$, 1.16$\times$, and 2.61$\times$ training speed up respectively (we define the moment when Fielding’s accuracy starts invariably exceeding the target accuracy as the point at which it achieves the target accuracy). On the highly heterogeneous and dynamic Open Images trace, Fielding achieves a significant accuracy improvement of 26%.
Although Auxo exhibits comparable or better performance than the global model baseline, the benefits are limited: accuracy improvement up to 0.9% and convergence speedup up to 1.05$\times$. These results are less impressive than those reported in the original paper on static datasets, suggesting a decline in clustering effectiveness. This degradation stems from re-clustering only the selected participants after each round, which neglects clients that have drifted but remain unselected. Additionally, this approach can lead to sudden accuracy drops on the biased Waymo Open dataset, where over 80% samples are cars, as selecting diverging clients can cause dramatic model weight changes. FlexCFL fails to improve model accuracy or convergence speed on all three real-world datasets. This is because FlexCFL migrates clients individually without adjusting the number of clusters. As we observe in Figure 1, such an approach will eventually make clusters as heterogeneous as a global set containing all clients.

### v2

@@ 2.2 System overview
Figure 3: Clustering overview of Fielding with label distribution as client representation. Clients send distribution vectors to the coordinator; the coordinator moves drifted clients to the closest cluster and triggers global re-clustering if any cluster center shifts by a distance larger than $\tau$.
2.3 Algorithmic framework
In this section, we describe our algorithm – presented in Algorithm 1 – in detail. Initially, Fielding clusters the clients using the $k$-means algorithm. The distance between clients is measured based on clients representations; in particular, for label-based clustering, we measure distance between data distributions, which is computed as the $\ell_{1}$-distance between their histograms: that is, assuming we have $l$ labels, the distance between clients with label distribution $p_{1},\ldots,p_{l}$ and $q_{1},\ldots,q_{l}$ is defined as $\sum_{i=1}^{l}|p_{i}-q_{i}|$.
We assume that the algorithm is executed for $T$ global iterations, and during each iteration, the label distribution of each client might change, for instance by adding or removing data points. The first step of Fielding is to handle data drift (Algorithm 2). Our goal is, by the end of each iteration, to maintain a good model for each cluster, with respect to its clients’ data. In this step, each drifted client is assigned to the closest cluster. If this causes a significant change in cluster centers – namely, if there exists a cluster moving by more than $\theta/3$, where $\theta$ is the average distance between cluster centers – we recluster clients from scratch. This condition avoids frequent reclustering, which might require excessive resources and can adversely change the clients’ losses. Importantly, we recluster all the clients, not just available clients, which, as we show in Appendix B, improves the accuracy.
For the global reclustering, we choose the number of clusters that gives the highest silhouette score. After reclustering, we compute new cluster models: for each client $i$, let $\mathbf{x}_{i}$ be its old cluster model, and then for each new cluster $C_{k}$, we define its model as an average of $\mathbf{x}_{i}$ for $i\in C_{k}$ (i.e., we average the old client models). After the clusters are computed, we train the cluster models for $R$ rounds. In each round, we sample clients from the cluster and run $L$ local iterations of gradient descent. After local iterations, we set the cluster model as the average of the models of all participating clients.
@@ 2.3 Algorithmic framework
1 Partition clients into clusters using $k$-means clusters based on client representations
@@ 2 Initialize cluster models $\mathbf{c}^{(1)},\ldots,\mathbf{c}^{(K)}$
4       Handle data drift using Algorithm 2
@@ 15                  $\mathbf{c}^{(k)}\leftarrow\operatorname*{avg}_{i\in S}\mathbf{x}_{i}$
Algorithm 1 CFL with data drift
@@ 7             Let $\mathbf{x}_{i}$ be the model corresponding to the $i$’th client’s cluster
8      Let $C_{1},\ldots,C_{K}$ be $k$-means clusters based on client representations
@@ 10             $\mathbf{c}^{{(k)}}\leftarrow\operatorname*{avg}_{i\in C_{k}}\mathbf{x}_{i}$
Algorithm 2 Handling data drift
1 Define the distance between clients based on their representations
@@ 2 Choose $K$ with the largest silhouette score
2.4 Convergence
In Appendix A, we analyze the performance of simplified version of our framework by bounding the average of clients’ loss functions at every iteration. The major challenge in the analysis is bounding the adverse effects of data drifts and reclustering. Data drifts change client datasets, hence changing their local objective functions. On the other hand, while reclustering is useful in the long run, its immediate effect on the objective can potentially be negative (see Figure 2(b)) due to the mixing of models from different clusters.
We make the following assumptions. [^1] See Appendix A for the precise statements and the discussion of the assumptions First, we assume that the distance between client representations – for example label distributions – translates into the difference between their objective functions. Second, we assume that the effect of data drift is bounded: that is, the representation $r_{i}$ of client $i$ changes by at most $\delta$ for each data drift. Finally, we assume that clients are clusterable; that is, there exists $K$ clusters so that representations of clients within each cluster are similar.
@@ 2.4 Convergence
Each $f^{(i)}_{t}$ – the local objective for client $i$ at iteration $t$ – is $L$-smooth and satisfies $\mu$-Polyak-Łojasiewicz condition, and we have access to a stochastic oracle with variance $\sigma^{2}$. There exists clustering such that representations inside each cluster are $\Delta$-close, the data drift changes representations by at most $\delta$, and the ratio between the difference between the objective function values and the distance between representations is at most $\theta$.
Under these assumptions, data drifts can affect each client representation – and hence the objective we optimize – by at most a fixed value. Moreover, by the clustering assumption, clients within each cluster have similar objectives after the reclustering, allowing us to bound the increase in the loss function due to reclustering. In Appendix A, we show the following result.
Theorem 1.
@@
$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$ $\leq(1-\eta\mu)^{tR}\left(f_{0}(\mathbf{x}_{0})-f_{0}(\mathbf{x}^{*})\right)$
$\quad+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
Intuitively, at every iteration, we provide a “regret” bound, comparing the loss at each cluster with the best loss we could have at each cluster. The bound has two terms: an exponentially decaying term corresponding to the initial loss, and a non-vanishing term corresponding to stochastic noise and the loss due to clustering and data drift. Ultimately, at early stages, the first term dominates, and we observe rapid improvement in objective value. On the other hand, at later stages, the first term vanishes, and the behavior is dictated by the stochastic and sampling noise, and severity of data drifts.
3 Evaluation
We evaluate Fielding prototype on four FL tasks with label distribution vectors as client representations (see Appendix E for results with other representations). End-to-end results highlight that Fielding improves final test accuracy by 1.9%-5.9% and is more stable than prior CFL methods. It works well with complementary FL optimizations, remains robust against malicious clients reporting corrupted distribution vectors, and still accelerates model convergence under low heterogeneity.
@@ 3 Evaluation
Environment. We emulate large-scale FL training with two GPU servers, each with two NVIDIA A100 GPUs (80 GB memory) and two AMD EPYC 7313 16-core CPUs. We use FedScale’s device datasets for realistic device computing and network capacity profiles while following the standardized training setup in the original paper (Lai et al., 2022). Per-device computation time is estimated from device computing speed, batch size, and number of local iterations; communication time is estimated based on the volume of downloaded and uploaded data and network bandwidth.
@@
Datasets and models. We use a satellite image dataset (FMoW (Christie et al., 2018)) and two video datasets (Waymo Open (Sun et al., 2020) and Cityscapes (Cordts et al., 2016)) that exhibit natural data drift. For FMoW, we keep images taken after January 1, 2015 and create one client per UTM zone. Two training rounds correspond to one day. For Waymo Open and Cityscapes, we use video segments pre-processed by Ekya (Bhardwaj et al., 2022). We create one client per camera per Waymo Open segment (212 clients in total), and one client per 100 consecutive frames per Cityscapes segment (217 clients in total). As these two datasets have video frame IDs within segments but lack global timestamps, we sort samples by frame ID and split them into 10 intervals. We stream in one data interval every 30 rounds with Cityscapes and every 20 rounds with Waymo Open. We train ResNet-18 (He et al., 2016) on Cityscapes and FMoW, and VisionTransformer-B16 (Dosovitskiy et al., 2021) on Waymo Open.
@@
To evaluate Fielding under highly dynamic drift, we construct a synthetic trace using the FedScale Open Images (Kuznetsova et al., 2020) benchmark. We find the top 100 most frequent classes and retain clients with samples from at least 10 of these 100 classes (5078 clients in total). Each client randomly partitions local data labels into 10 buckets, with each bucket then containing all samples of its labels. We stream in one bucket every 50 rounds and use ShuffleNet v2 (Ma et al., 2018) for this task. For all four datasets, clients start with 100 rounds worth of data and retain samples from the most recent 100 rounds. Dataset and training configurations are summarized in Table 1 and Table 2.
@@
Metrics. Our main evaluation metrics are Time-to-Accuracy (TTA) and final test accuracy. We define TTA as the training time required to achieve the target accuracy–the average client test accuracy when training a single global model (black dashed lines in Figure 4). We report the final average test accuracy across diverse settings to demonstrate the sensitivity and robustness of Fielding.
Figure 4: Time to accuracy (TTA) comparison over four tasks.
3.1 End-to-end training performance
Figure 4 shows that Fielding improves average client test accuracy by 5.9%, 1.9%, and 3.4% in FMoW, Cityscapes, and Waymo Open, respectively. These improvements are significant, given that baseline test accuracies range from 45% to 85%, and are comparable to the accuracy gains prior CFL works achieve on static data. When considering the global model’s final test accuracy as the target accuracy, Fielding offers a 1.27$\times$, 1.16$\times$, and 2.23$\times$ training speed-up, respectively (We define the point at which Fielding’s accuracy consistently surpasses the target accuracy as the moment it achieves the target accuracy). On the highly heterogeneous and dynamic Open Images trace, Fielding boosts accuracy by 17.9-26.1% in the final 100 rounds.
While Auxo performs comparably or slightly better than the global model baseline, the benefits are limited: up to 0.9% accuracy gain and up to 1.05$\times$ convergence speedup. These results fall short of those reported in the original paper on static datasets, suggesting a decline in clustering effectiveness. This degradation stems from re-clustering only the selected participants after each round, ignoring drifted but unselected clients. This approach also leads to sudden accuracy drops on the biased Waymo Open dataset, where over 80% samples are cars, due to sudden model shifts when divergent clients are selected. FlexCFL shows no improvement in accuracy or convergence across all real-world datasets, as it migrates clients individually without adjusting the number of clusters. As Figure 1 shows, this approach leads to cluster heterogeneity approaching that of the global client set.
FedDrift Comparison. FedDrift (Jothimurugesan et al., 2023) is another clustered FL method designed to handle drift. However, it is impractical for large-scale settings, as it (1) requires every client to train every cluster model in each round and (2) assumes all clients remain online and available for training. Therefore, we construct a small-scale setting with the FMoW dataset and 50 clients (the number of clients we selected per round in our original experiment). Figure 10 shows that Fielding achieves 2.5% higher final accuracy and reaches FedDrift’s final accuracy 2.01$\times$ faster.
3.2 Compatibility with other optimizations
To demonstrate that Fielding is agnostic to client selection and model aggregation strategies, we evaluate Fielding’s performance on FMoW together with various complementary optimizations. In Figure 5 and Figure 6, we use the algorithm name alone to denote the baseline where we train one global model and use "+ Fielding " to denote running the algorithm atop Fielding.
We use the Oort selection algorithm (Lai et al., 2021) and a distance-based algorithm prioritizing clients closer to the distribution center as client selection examples. As shown in Figure 5, Fielding improves the final average test accuracy by 6.0% and 4.5% and reach the target accuracy 2.62×2.62$\times$2.62 × and 1.17×1.17$\times$1.17 × faster. Note that Oort selection actively incorporates clients otherwise filtered out due to long response time, leading to a longer average round time. Figure 6 shows that Fielding works well with aggregation algorithm FedYogi (Reddi et al., 2021) and q-FedAvg (Li et al., 2020c). Fielding improves the final accuracy by 5.9% and 4.9% while giving a 1.34×1.34$\times$1.34 × and 1.25×1.25$\times$1.25 × speedup respectively.
3.3 Robustness and sensitivity analysis